SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Nov 14th, 2006)

Attendees: Debbie Bard, Rainer Bartoldus, Al Eisner, Fang Fang, Selina Li, Rahmat Rahmat, Su Dong, Mark Tibbetts.

Operations

  • The old TSFX crate in the EH was removed on Friday as requested, because the space is needed for the second gas bubbler. Debbie will speak to Ray about scheduling the re-cabling of the fibers. Also, he monitoring cable chain was broken by removing that crate - the DCZ isn't being monitored any more, so either the cable can be re-routed directly to the new rack or a junction connector can be used. Speak to Ray about this as well.

EMT

  • The modifications made to EmtSim package are summerised here. Mark will run Moose to compare events before and after the modifications were made, for validation.

Online

  • Fang Fang found the problem wiht the test-stand was a faulty cable from the ROM to the TSF board. Swapping the cable solved the problem. Still problems with the Dr. test set-up, Fang Fang will talk to Rainer about using the database to fix this.
  • A different test-stand problem came up - an FPGA didn't boot when first turned on. It works now, but that's never happened before....

L3

  • The first stage of the Trigger analysis survey is effrectively done - see here. The next stage involves deployment and validation, and that's underway.

Data Quality and Monitoring

  • The memory leak problems mentioned last week has been solved and the next production release will contain the code changes.

AOB

  • The LST installation has vbeen completed, so we need to check with the IFR IFT expert that the neccessary changes have been/will be made to the IFT boards.
  • Any change of IFT object should be included in the IFT documentaion.

This page is created by Debbie Bard (djbard@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.